OpenStudy (anonymous):

Which sequence follows the rule 7n-2, where n represents the position of a term in the sequence?

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@nincompoop ill give it a shot at tagging you.

OpenStudy (anonymous):

@mathgirl73

OpenStudy (anonymous):

please

OpenStudy (anonymous):

im going to fan you

OpenStudy (anonymous):

Okay?

OpenStudy (anonymous):

@nubeer

OpenStudy (nubeer):

Arithmetic prgression maybe

OpenStudy (anonymous):

okay so? can you help? medals are awarded.

OpenStudy (anonymous):

see?

OpenStudy (nubeer):

lol medals really doesnt mattter for me.. if i can help i would

OpenStudy (anonymous):

so. u cant. ok.

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

@zepp

OpenStudy (zepp):

So, where are the sequences

OpenStudy (anonymous):

the seauences are not given. uve to make one up, i think

OpenStudy (anonymous):

o no wait.

OpenStudy (anonymous):

a. 10,12.14,16

OpenStudy (anonymous):

b. 14,21,28,35

OpenStudy (anonymous):

c. 5,12,19,26

OpenStudy (anonymous):

d. 5,15,20,25

OpenStudy (anonymous):

There we go. Sorry *facepalm*

OpenStudy (zepp):

Try to plug 1,2,3,4 in your rule, For n=1, 7n-2 = 7-2=5 For n=2, 7n-2 = 14-2 = 12 etc, and you'll find out that the first number of this sequence, beginning with a 5, goes up of 7 everytime, so 5,12,19,26,33,40 and so on.

OpenStudy (anonymous):

I got the answer.

OpenStudy (anonymous):

it was c